Consider the following investment projects: 1 AC DE 0-1500-2000-2100-1000 1 3000 1560 2800 3600 2-1500 944 -200 -5700 3 15 0 marr 13%

the MARR is 12%
Answer #1
a Number of possible IRR =Number of sign changes
Project Number of possible IRR
A 3
C 1
D 2
E 3
b Simple Investments are investments with only one cahnge in sign in the net cash flow
Non Simple Investments are investments with more than one cahnge in sign in the net cash flow
Project
A Non-Simple
C Simple
D Non-Simple
E Non-Simple
Present Value (PV)of Cash Flow=(Cash Flow)/((1+i)^N)
i=discount rate , N=Year of cash Flow
NPV =Sum of PV
Project A
N CF PV1=CF/(1^N) PV2=CF/(1.1^N) PV3 PV4 PV5 PV6 PV7 PV8 PV9 PV10 PV11 PV12 PV13
Year Cash Flow PV at 0% PV at 10% PV at 20% PV at 30% PV at 40% PV at 50% PV at 60% PV at 70% PV at 80% PV at 90% PV at 100% PV at110% PV at 120%
0 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500 -1500
1 3000 3000 2727.272727 2500 2307.692 2142.857143 2000 1875 1764.706 1666.667 1578.947 1500 1428.5714 1363.636
2 -1500 -1500 -1239.669421 -1041.67 -887.574 -765.3061224 -666.667 -585.938 -519.031 -462.963 -415.512 -375 -340.13605 -309.917
3 15 15 11.26972201 8.680556 6.827492 5.466472303 4.444444 3.662109 3.053124 2.572016 2.186908 1.875 1.6196955 1.408715
Discount Rate 0% 10% 20.0% 30.0% 40.0% 50.0% 60.0% 70.0% 80.0% 90.0% 100.0% 110.0% 120.0%
NPV 15.00 -1.13 -32.99 -73.05 -116.98 -162.22 -207.28 -251.27 -293.72 -334.38 -373.13 -409.94 -444.87
SUM of PV
IRR 10%
Project C
Project C N PV1=CF/(1AN) PV2=CF/(1.1^N) PV3 PV6 PV7 PV13 CF PV4 PV5 PV8 PV9 PV10 PV11 PV12 Year Cash Flow PV at 10% PV at 20%Project D PV1-CF/(1AN) PV2=CF/(1.1^N) PV3 PV4 PV12 CF PV5 PV6 PV7 PV8 PV9 PV10 PV11 PV13 Cash Flow PV at 0% PV at 10% PV at 2NPV PROFIE -PROJECT A 100.00 O 0.00 140% O % 40% 60% 80% 100% 120% -100.00 -200.00 O -300.00 -400.00 -500.00 ONPV PROFILE_PROJECT C 0.00 40.0 % 0.0% 20.0% 60.0% 80.0% 100.0% 120.0% 140.0% -200.00 400.00 -600.00 -800.00 -1000.00 -1200.0NPV PROFILE_PROJECT D 600.00 400.00 200.00 0.00 O% 100% 120% 20% 40% 60% 80% 140% -200.00 -400.00 -600.00 -800.00 -1000.00NPV PROFILE_PROJECT E 600.00 400.00 200.00 0.00 O% 60% 20% 40% 80% 100% 120% 140% -200.00 -400.00 -600.00 -800.00 -1000.00fox Е32 X IRR(D25:D28) D Е F G Н к L M N 23 N PV1=CF/(1AN) PV2=CF/(1.1^N) PV3 PV PV6 PV7 CF PV4 PV5 PV8 PV9 PV10 PV at 50% PVfx =IRR(D36:D39) E42 C D E G К L M O 33 Project C PV1=CF/(1AN) PV2=CF/(1.1^N) PV3 34 N PV1 PV7 PV11 CF PV4 PV5 PV6 PV8 PV9 PVfc =IRR(D60:D63) Еб7 A Е G Н -1000 48 0 -1000 -1000 -1000 -1000 3600 3600 25 49 1 3272.727273 3000 2769.231 -4710.743802 -395